# Two Tales of Single-Phase Contrastive Hebbian Learning

### v2

@@ 4.3 An Adversarial Relaxed Optimal Value Reformulation
Let $0<\beta\leq\beta^{\prime}$ and $\alpha,\alpha^{\prime}\in[0,1]$ with $\alpha\leq\alpha^{\prime}$. Then the following holds: (a) $J_{\alpha,\beta}^{\text{AROVR}}(\theta)\leq J_{\alpha^{\prime},\beta}^{\text{AROVR}}(\theta)$ and (b) $\beta J_{\alpha,\beta}^{\text{AROVR}}(\theta)\leq\beta^{\prime}J_{\alpha,\beta^{\prime}}^{\text{AROVR}}(\theta)$.
Original EP ($\alpha=1$) and centered EP [26, 42] ($\alpha=1/2$) are two particular instances of $J_{\alpha,\beta}^{\text{AROVR}}$. Hence, Prop. 4.1 raises the question of how much the better gradient estimate and how much the stricter loss $J_{1/2,\beta}^{\text{AROVR}}$ contribute to the advantages of centered EP.
@@ Appendix B CNN experimental details
Let $0<\beta\leq\beta^{\prime}$ and $\alpha,\alpha^{\prime}\in[0,1]$ with $\alpha\leq\alpha^{\prime}$. Then the following holds: (a) $J_{\alpha,\beta}^{\text{EP}}(\theta)\leq J_{\alpha^{\prime},\beta}^{\text{EP}}(\theta)$ and (b) $\beta J_{\alpha,\beta}^{\text{EP}}(\theta)\leq\beta^{\prime}J_{\alpha,\beta^{\prime}}^{\text{EP}}(\theta)$.
@@
Claim (a) In the following we absorb $1/\beta$ into $E$ and omit the dependency of $E$ on $\theta$ for notational brevity. $\alpha\leq\alpha^{\prime}$ implies $\bar{\alpha}\leq\bar{\alpha}^{\prime}\leq 0\leq\alpha\leq\alpha^{\prime}$. Further, we define
$\begin{split}s_{\alpha}&:=\arg\min_{s}\alpha\ell(s)+E(u)\qquad s_{-\bar{\alpha}}:=\arg\min_{s}-\bar{\alpha}\ell(s)+E(s)\\ s_{\alpha^{\prime}}&:=\arg\min_{s}\alpha^{\prime}\ell(s)+E(u)\qquad s_{-\bar{\alpha}^{\prime}}:=\arg\min_{s}-\bar{\alpha}^{\prime}\ell(s)+E(s).\end{split}$ (33)
We identify $f(s)=\ell(s)$ and $g(s)=E(s,\theta)$ and apply the lemma to deduce that
@@
$\begin{split}J_{\alpha,\beta}^{\text{EP}}(\theta)&=\alpha\ell(s_{\alpha})+E(s_{\alpha})+\bar{\alpha}\ell(s_{-\bar{\alpha}})-E(s_{-\bar{\alpha}})\\ &\leq\alpha\ell(s_{\alpha^{\prime}})+E(s_{\alpha^{\prime}})+\bar{\alpha}\ell(s_{-\bar{\alpha}})-E(s_{-\bar{\alpha}})\\ &=\alpha^{\prime}\ell(s_{\alpha^{\prime}})+E(s_{\alpha^{\prime}})+(\alpha-\alpha^{\prime})\ell(s_{\alpha^{\prime}})+\bar{\alpha}\ell(s_{-\bar{\alpha}})-E(s_{-\bar{\alpha}})\\ &\leq\alpha^{\prime}\ell(s_{\alpha^{\prime}})+E(s_{\alpha^{\prime}},)+(\alpha-\alpha^{\prime})\ell(s_{-\bar{\alpha}})+\bar{\alpha}\ell(s_{-\bar{\alpha}})-E(s_{-\bar{\alpha}})\\ &=\alpha^{\prime}\ell(s_{\alpha^{\prime}})+E(s_{\alpha^{\prime}})+\bar{\alpha}^{\prime}\ell(s_{-\bar{\alpha}})-E(s_{-\bar{\alpha}})\\ &\leq\alpha^{\prime}\ell(s_{\alpha^{\prime}})+E(s_{\alpha^{\prime}})+\bar{\alpha}^{\prime}\ell(s_{-\bar{\alpha}^{\prime}})-E(s_{-\bar{\alpha}^{\prime}})\\ &=J_{\alpha^{\prime},\beta}^{\text{EP}}(\theta),\end{split}$ (35)
where we used the definition of $s_{\alpha}$ in the first inequality, $\ell(s_{\alpha^{\prime}})\leq\ell(s_{-\bar{\alpha}})$ (together with $\alpha-\alpha^{\prime}\geq 0$) for the second one, and finally the definition of $s_{-\bar{\alpha}^{\prime}}$ to obtain the last inequality.
@@
$\begin{split}\beta J_{\alpha,\beta}^{\text{EP}}(\theta)&=\gamma\ell(s_{\gamma})+E(s_{\gamma})+\bar{\gamma}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &\leq\gamma\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\bar{\gamma}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &=\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+(\gamma-\gamma^{\prime})\ell(s_{\gamma^{\prime}})+\bar{\gamma}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &\leq\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+(\gamma-\gamma^{\prime})\ell(s_{-\bar{\gamma}})+\bar{\gamma}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &=\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\underbrace{(\gamma-\gamma^{\prime}+\bar{\gamma})}_{=\bar{\gamma}^{\prime}}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &\leq\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\bar{\gamma}^{\prime}\ell(s_{-\bar{\gamma}^{\prime}})-E(s_{-\bar{\gamma}^{\prime}})\\ &=\beta^{\prime}J_{\alpha,\beta^{\prime}}^{\text{EP}}(\theta).\end{split}$ (40)
We applied the definition of $s_{\gamma}$ (first inequality), $\ell(s_{\gamma^{\prime}})\leq\ell(s_{-\bar{\gamma}})$ (2nd inequality) and the definition of $s_{\bar{\gamma}^{\prime}}$. ∎
@@
Let $0<\beta^{\prime}<\beta$ and $\alpha^{\prime}$ such that $\bar{\alpha}\beta=\bar{\alpha}^{\prime}\beta^{\prime}$. Then $J_{\alpha,\beta}^{\text{EP}}(\theta)\leq J_{\alpha^{\prime},\beta^{\prime}}^{\text{EP}}(\theta)$.
@@
$\tilde{U}_{-\gamma}(s):=-\gamma\ell(s)+U(s)$ (41)
@@
$\begin{split}s_{\alpha\beta}&=\arg\min_{s}\alpha\ell(s)+\tfrac{1}{\beta}U(s)\\ s_{\alpha^{\prime}\beta^{\prime}}&=\arg\min_{s}\alpha^{\prime}\ell(s)+\tfrac{1}{\beta^{\prime}}U(s)\\ s_{-\gamma}&=\arg\min_{s}-\bar{\alpha}\ell(s)+\tfrac{1}{\beta}U(s)=\arg\min_{s}\tilde{U}_{-\gamma}(s).\end{split}$ (42)
@@
$\begin{split}J_{\alpha,\beta}^{\text{EP}}(\theta)&=\alpha\ell(s_{\alpha\beta})+\tfrac{1}{\beta}U(s_{\alpha\beta})+\bar{\alpha}\ell(s_{-\gamma})-\tfrac{1}{\beta}U(s_{-\gamma})\\ &=\ell(s_{\alpha\beta})+\tfrac{1}{\beta}\tilde{U}_{-\gamma}(s_{\alpha\beta})-\tfrac{1}{\beta}\tilde{U}_{-\gamma}(s_{-\gamma})\\ &\leq\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta}\tilde{U}_{-\gamma}(s_{\alpha^{\prime}\beta^{\prime}})-\tfrac{1}{\beta}\tilde{U}_{-\gamma}(s_{-\gamma})\\ &=\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta}\underbrace{\left(\tilde{U}_{-\gamma}(s_{\alpha^{\prime}\beta^{\prime}})-\tilde{U}_{-\gamma}(s_{-\gamma})\right)}_{\geq 0}\\ &\leq\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta^{\prime}}\left(\tilde{U}_{-\gamma}(s_{\alpha^{\prime}\beta^{\prime}})-\tilde{U}_{-\gamma}(s_{-\gamma})\right)\\ &=\alpha^{\prime}\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta^{\prime}}U(s_{\alpha^{\prime}\beta^{\prime}})+\bar{\alpha}^{\prime}\ell(s_{-\gamma})-\tfrac{1}{\beta^{\prime}}U(s_{-\gamma})=J_{\alpha^{\prime},\beta^{\prime}}^{\text{EP}}(\theta),\end{split}$ (43)

### v3

@@ 4.3 An Adversarial Relaxed Optimal Value Reformulation
Let $0<\beta\leq\beta^{\prime}$ and $\alpha,\alpha^{\prime}\in[0,1]$ with $\alpha\leq\alpha^{\prime}$. Then the following holds: (a) $J_{\alpha^{\prime},\beta}^{\text{AROVR}}(\theta)\leq J_{\alpha,\beta}^{\text{AROVR}}(\theta)$ and (b) $\beta J_{\alpha,\beta}^{\text{AROVR}}(\theta)\leq\beta^{\prime}J_{\alpha,\beta^{\prime}}^{\text{AROVR}}(\theta)$.
Original EP ($\alpha=1$) and centered EP [26, 42] ($\alpha=1/2$) are two particular instances of $J_{\alpha,\beta}^{\text{AROVR}}$. Hence, Prop. 4.1 raises the question of how much the better gradient estimate and how much the stricter loss $J_{1/2,\beta}^{\text{AROVR}}$ contribute to the advantages of centered EP over standard EP.
@@ Appendix B CNN experimental details
Let $0<\beta\leq\beta^{\prime}$ and $\alpha,\alpha^{\prime}\in[0,1]$ with $\alpha\leq\alpha^{\prime}$. Then the following holds: (a) $J_{\alpha^{\prime},\beta}^{\text{AROVR}}(\theta)\leq J_{\alpha,\beta}^{\text{AROVR}}(\theta)$ and (b) $\beta J_{\alpha,\beta}^{\text{AROVR}}(\theta)\leq\beta^{\prime}J_{\alpha,\beta^{\prime}}^{\text{AROVR}}(\theta)$.
@@
Claim (a) In the following we absorb $1/\beta$ into $E$ and omit the dependency of $E$ on $\theta$ for notational brevity. $\alpha\leq\alpha^{\prime}$ implies $-\bar{\alpha}\leq-\bar{\alpha}^{\prime}\leq 0\leq\alpha\leq\alpha^{\prime}$. Further, we define
$\begin{split}s_{\alpha}&:=\arg\min_{s}\alpha\ell(s)+E(s)\qquad s_{-\bar{\alpha}}:=\arg\min_{s}-\bar{\alpha}\ell(s)+E(s)\\ s_{\alpha^{\prime}}&:=\arg\min_{s}\alpha^{\prime}\ell(s)+E(s)\qquad s_{-\bar{\alpha}^{\prime}}:=\arg\min_{s}-\bar{\alpha}^{\prime}\ell(s)+E(s).\end{split}$ (33)
We identify $f(s)=\ell(s)$ and $g(s)=E(s,\theta)$ and apply the lemma repeatedly to deduce that
@@
$\begin{split}J_{\alpha^{\prime},\beta}^{\text{AROVR}}(\theta)&=\underbrace{\alpha^{\prime}\ell(s_{\alpha^{\prime}})+E(s_{\alpha^{\prime}})}_{\leq\alpha^{\prime}\ell(s_{\alpha})+E(s_{\alpha})}+\bar{\alpha}^{\prime}\ell(s_{-\bar{\alpha}^{\prime}})-E(s_{-\bar{\alpha}^{\prime}})\\ &\leq\alpha^{\prime}\ell(s_{\alpha})+E(s_{\alpha})+\bar{\alpha}^{\prime}\ell(s_{-\bar{\alpha}^{\prime}})-E(s_{-\bar{\alpha}^{\prime}})\\ &=\alpha\ell(s_{\alpha})+E(s_{\alpha})+\bar{\alpha}\ell(s_{-\bar{\alpha}^{\prime}})-E(s_{-\bar{\alpha}^{\prime}})+(\alpha^{\prime}-\alpha)\ell(s_{\alpha})+(\bar{\alpha}^{\prime}-\bar{\alpha})\ell(s_{-\bar{\alpha}^{\prime}})\\ &=\alpha\ell(s_{\alpha})+E(s_{\alpha})+\bar{\alpha}\ell(s_{-\bar{\alpha}^{\prime}})-E(s_{-\bar{\alpha}^{\prime}})+(\alpha^{\prime}-\alpha)\underbrace{\big{(}\ell(s_{\alpha})-\ell(s_{-\bar{\alpha}^{\prime}})\big{)}}_{\leq 0}\\ &\leq\alpha\ell(s_{\alpha})+E(s_{\alpha})+\underbrace{\bar{\alpha}\ell(s_{-\bar{\alpha}^{\prime}})-E(s_{-\bar{\alpha}^{\prime}})}_{\leq\bar{\alpha}\ell(s_{-\bar{\alpha}})-E(s_{-\bar{\alpha}})}\\ &\leq\alpha\ell(s_{\alpha})+E(s_{\alpha})+\bar{\alpha}\ell(s_{-\bar{\alpha}})-E(s_{-\bar{\alpha}})\\ &=J_{\alpha,\beta}^{\text{AROVR}}(\theta),\end{split}$ (35)
where we used the definition of $s_{\alpha}$ in the first inequality, $\ell(s_{\alpha})\leq\ell(s_{-\bar{\alpha}^{\prime}})$ (together with $\alpha^{\prime}-\alpha\geq 0$) for the second one, and finally the definition of $s_{-\bar{\alpha}}$ to obtain the last inequality.
@@
$\begin{split}\beta J_{\alpha,\beta}^{\text{AROVR}}(\theta)&=\gamma\ell(s_{\gamma})+E(s_{\gamma})+\bar{\gamma}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &\leq\gamma\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\bar{\gamma}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &=\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\bar{\gamma}^{\prime}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})+(\gamma-\gamma^{\prime})\ell(s_{\gamma^{\prime}})+(\bar{\gamma}-\bar{\gamma}^{\prime})\ell(s_{-\bar{\gamma}})\\ &=\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\bar{\gamma}^{\prime}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})+(\gamma-\gamma^{\prime})\underbrace{\big{(}\ell(s_{\gamma^{\prime}})-\ell(s_{-\bar{\gamma}})\big{)}}_{\leq 0}\\ &\leq\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\bar{\gamma}^{\prime}\ell(s_{-\bar{\gamma}})-E(s_{-\bar{\gamma}})\\ &\leq\gamma^{\prime}\ell(s_{\gamma^{\prime}})+E(s_{\gamma^{\prime}})+\bar{\gamma}^{\prime}\ell(s_{-\bar{\gamma}^{\prime}})-E(s_{-\bar{\gamma}^{\prime}})\\ &=\beta^{\prime}J_{\alpha,\beta^{\prime}}^{\text{AROVR}}(\theta).\end{split}$ (40)
We applied the definition of $s_{\gamma}$ (first inequality), $\ell(s_{\gamma^{\prime}})\leq\ell(s_{-\bar{\gamma}})$ (2nd inequality) and the definition of $s_{-\bar{\gamma}^{\prime}}$. ∎
@@
$\tilde{E}_{-\gamma}(s):=-\gamma\ell(s)+E(s)$ (41)
@@
$\begin{split}s_{\alpha\beta}&=\arg\min_{s}\alpha\ell(s)+\tfrac{1}{\beta}E(s)\\ s_{\alpha^{\prime}\beta^{\prime}}&=\arg\min_{s}\alpha^{\prime}\ell(s)+\tfrac{1}{\beta^{\prime}}E(s)\\ s_{-\gamma}&=\arg\min_{s}-\bar{\alpha}\ell(s)+\tfrac{1}{\beta}E(s)=\arg\min_{s}\tilde{E}_{-\gamma}(s).\end{split}$ (42)
@@
$\begin{split}J_{\alpha,\beta}^{\text{AROVR}}(\theta)&=\alpha\ell(s_{\alpha\beta})+\tfrac{1}{\beta}E(s_{\alpha\beta})+\bar{\alpha}\ell(s_{-\gamma})-\tfrac{1}{\beta}E(s_{-\gamma})\\ &\leq\alpha\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta}E(s_{\alpha\beta})+\bar{\alpha}\ell(s_{-\gamma})-\tfrac{1}{\beta}E(s_{-\gamma})\\ &=\ell(s_{\alpha^{\prime}\beta^{\prime}})-\bar{\alpha}\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta}E(s_{\alpha\beta})+\bar{\alpha}\ell(s_{-\gamma})-\tfrac{1}{\beta}E(s_{-\gamma})\\ &=\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta}\tilde{E}_{-\gamma}(s_{\alpha^{\prime}\beta^{\prime}})-\tfrac{1}{\beta}\tilde{E}_{-\gamma}(s_{-\gamma})\\ &=\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta}\underbrace{\left(\tilde{E}_{-\gamma}(s_{\alpha^{\prime}\beta^{\prime}})-\tilde{E}_{-\gamma}(s_{-\gamma})\right)}_{\geq 0}\\ &\leq\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta^{\prime}}\left(\tilde{E}_{-\gamma}(s_{\alpha^{\prime}\beta^{\prime}})-\tilde{E}_{-\gamma}(s_{-\gamma})\right)\\ &=\alpha^{\prime}\ell(s_{\alpha^{\prime}\beta^{\prime}})+\tfrac{1}{\beta^{\prime}}E(s_{\alpha^{\prime}\beta^{\prime}})+\bar{\alpha}^{\prime}\ell(s_{-\gamma})-\tfrac{1}{\beta^{\prime}}E(s_{-\gamma})=J_{\alpha^{\prime},\beta^{\prime}}^{\text{AROVR}}(\theta),\end{split}$ (43)